# Universal Entanglement Revival of Topological Origin

### v1

@@ Appendix B Entropy density
We also consider the average entropy $s_{\gamma}(\hat{\rho})=S_{\gamma}(\hat{\rho})/n$, where $S_{\gamma}(\hat{\rho})$ is defined in Eq. (23) and $n=(L-1)/2$ is the number of two-site unit cells. These quantities can be easily evaluated within the semiclassical approximation of $\hat{\rho}$, given by Eqs. (14) and (15). In particular, the von Neumann entropy density $s_{\text{vN}}(\rho)$ is obtained as follows:
@@
and the Rényi entropy density of order 2 reads:
@@
The qualitative time dependence of these quantities differs from the logarithmic negativity, since the system is in a pure state at $t=0$ and $t\to\infty$ (when the entropy is zero) and there is no revival behavior. Instead, the entropy density attains its maximum value $s_{\mathrm{vN}}=s_{2}=1$ at $\Gamma t=2\ln 2$. The increase of the entropy at intermediate times in in qualitative agreement with the reduction of purity and the occurrence of a minimum in the logarithmic negativity. However, in contrast to the entanglement, the maximum entropy always occurs at $\Gamma t=2\ln 2$, independent of the coupling strengths $g_{1,2}$. This difference is due to the fact that the probabilities of the eigenstates, given in Eq. (15), do not depend on $g_{1,2}$. Instead, the entanglement is obviously affected by the specific eigenstates which diagonalize $\hat{\rho}$, thus the position of the minimum depends the coupling ratio $g_{1}/g_{2}$. Despite these remarks, we note that both Eqs. (24) and (25) are independent on $L$, i.e., they display a universal behavior, demonstrating that local quantum correlations are robust with respect to the overall system size. These results suggest that the robustness of local quantum correlations is likely due to the topological properties of the SSH model, contributing to the stability of the entanglement revival under different conditions.

### v2

@@ Appendix B Entropy density
[…]
@@
and the Rényi entropy density of order 2 reads as
@@
The qualitative time dependence of these quantities features dynamical behavior distinct from that of entanglement measures such as logarithmic negativity. Since the system is in a pure state at $t=0$ and $t\to\infty$, they vanish at $t=0$ and $t=\infty$. Instead, the entropy density attains its maximum value $s_{\mathrm{vN}}=s_{2}=1$ at $\Gamma t=2\ln 2$. The increase of the entropy at intermediate times is in qualitative agreement with the reduction of purity and the occurrence of a minimum in the logarithmic negativity. Interestingly, in contrast to entanglement, the maximum entropy always occurs at $\Gamma t=2\ln 2$, independent of the coupling strengths $g_{1,2}$. This difference is due to the fact that the probabilities of the eigenstates, given in Eq. (20), do not depend on $g_{1,2}$. Instead, the entanglement is obviously affected by the specific eigenstates of $\hat{\rho}$, thus the position of the minimum depends the coupling ratio $g_{1}/g_{2}$. Despite these differences, we note that both Eqs. (31) and (32) are independent of $L$, i.e., they also display a universal behavior, demonstrating that local quantum correlations are robust with respect to the overall system size. These results suggest that the robustness of local quantum correlations is likely due to the topological properties of the SSH model, contributing to the stability of the entanglement revival under different conditions.
Appendix C Spatially correlated noise
Figure A4: Entanglement dynamics under nearest-neighbor correlated decoherence, $\hat{L}_{i}=\hat{f}_{2i-1}+\hat{f}_{2i+1}$. The entanglement is shown for various system sizes, ranging from $L=7$ to $L=1001$, and under different coupling ratios: (a) $g_{1}/g_{2}=0.1$, (b) $g_{1}/g_{2}=0.5$, and (c) $g_{1}/g_{2}=0.9$. The parameter that we take: $g_{2}=1,\Gamma/g_{2}=0.1$
Figure A5: Entanglement dynamics under exponentially decaying correlated decoherence, $\hat{L}_{i}=\hat{f}_{2i-1}+e^{-1}(\hat{f}_{2i+1}+\hat{f}_{2i-3})+e^{-2}(\hat{f}_{2i+3}+\hat{f}_{2i-5})+\dots$. The entanglement is shown for various system sizes, ranging from $L=5$ to $L=1001$, and under different coupling ratios: (a) $g_{1}/g_{2}=0.1$, (b) $g_{1}/g_{2}=0.5$, and (c) $g_{1}/g_{2}=0.9$. The parameter that we take: $g_{2}=1,\Gamma/g_{2}=0.1$.
In the main text we assumed for simplicity a strictly local decoherence channel. In realistic platforms, however, spatially correlated noise may arise and influence the dynamics. Inter-qubit noise correlations have been experimentally identified in densely packed silicon spin qubits Yoneda et al. (2023), and their effects on qubit dynamics have been theoretically analyzed in both Markovian and non-Markovian regimes Zou et al. (2024). These studies underscore the importance of considering correlated noise when investigating entanglement under decoherence.
To test the robustness of our results, we consider a decay channel with spatial correlations, which nevertheless remain short-ranged. This distinction is crucial, as topological properties are robust against local fluctuations and disorder, but can break down under globally correlated perturbations. For definiteness, here we restrict ourselves to the fermionic case and examine two representative types of correlated Lindblad operators that preserve the chiral symmetry of the SSH model: (i) The nearest-neighbor symmetric combination $\hat{L}_{i}=\hat{f}_{2i-1}+\hat{f}_{2i+1}$, where decoherence acts collectively on nearest-neighbor odd sites; (ii) The exponentially decaying form $\hat{L}_{i}=\hat{f}_{2i-1}+e^{-1}(\hat{f}_{2i+1}+\hat{f}_{2i-3})+e^{-2}(\hat{f}_{2i+3}+\hat{f}_{2i-5})+\dots$, which introduces a broader spatial profile of correlated decoherence, mimicking decaying interactions often present in realistic environments.
As shown in Fig. A4 and A5, the entanglement dynamics under such spatially correlated decoherence remains qualitatively similar to that observed with local noise (cf. Fig. 3). The universal entanglement revival persists across various coupling ratios and becomes more pronounced in larger systems. We note that for $g_{1}/g_{2}=0.9$, the universal behavior is less visible in small systems [see Fig. A4 (c) and A5 (c)] since the system size must exceed the localization length of edge mode, which grows fast as $g_{1}/g_{2}$ approaches 1. All these results demonstrate that the universal entanglement revival persists for spatially correlated noise, as long as it preserves the chiral symmetry and the correlation length is not too long compared with the system size. In fact, as already noted, topological states are protected against local perturbations, but not global ones.
Appendix D Finite temperature
Figure A6: Entanglement dynamics, assuming that the small fraction $\delta=0.05$ of negative-energy modes with the highest energy is left unoccupied, while all deeper modes and the edge mode remain filled. Explicitly, the initial state is $|\psi_{g}\rangle=\hat{c}_{0}^{\dagger}\prod_{k=1}^{(L-1)/2-\delta(L-1)}\hat{c}_{-k}^{\dagger}\mathinner{\lvert{\textstyle\ }\rangle}$. We take $\Gamma/g_{2}=0.1$.
Figure A7: Entanglement dynamics including thermal activation processes, simulated by adding Lindblad operators $\hat{L}_{i}=\hat{f}_{2i-1}^{\dagger}$. The ratio between the excitation and decay rates is set to $\Gamma_{\uparrow}/\Gamma_{\downarrow}=0.5$, and $\Gamma_{\downarrow}/g_{2}=0.1$.
In realistic quantum platforms, finite temperature is unavoidable. It may affect the preparation of the initial state, as well as the subsequent dynamics. To examine whether the entanglement revival behavior persists under such conditions, we consider two main effects: (i) thermal excitations in the initial state, and (ii) thermal activation processes during the dynamics.
In the first case, we suppose that the temperature $T$ satisfies $T\ll g_{1,2}$, and that the chemical potential $\mu$ is close to the upper edge of the negative-energy band. From Eq. (4), the latter condition is easily written as $\mu\approx-g_{2}+g_{1}$, and is realized when $k\approx(L-1)/2$. Since the energy gap is $2(g_{2}-g_{1})$, the population of the positive-energy states remains negligible (at least, sufficiency far from the boundary of the topological phase). It follows that the main consequence of a finite temperature is the depopulation of energy states close to the chemical potential, while deeper negative-energy modes and the zero-energy edge mode remain occupied. To capture this effect, we consider the initial state $|\psi_{g}\rangle=\hat{c}_{0}^{\dagger}\prod_{k=1}^{(L-1)/2-\delta(L-1)}\hat{c}_{-k}^{\dagger}\,\mathinner{\lvert{\textstyle\ }\rangle}$, where $\delta$ indicates the small fraction of unoccupied negative-energy states. As shown in Fig. A6, the universal entanglement revival behavior exhibits excellent robustness.
Although the above result has been obtained by assuming, for simplicity, a pure initial state, we should also stress that the system undergoes a dramatic loss of purity in the course of the time evolution, due to the simultaneous decay of all occupied modes. Therefore, even for a thermally mixed initial state, we do not expect significant qualitative changes to the revival dynamics in the low-temperature regime $T\ll g_{1,2}$. The main difference will be a small decrease of purity (and entanglement) at the beginning of the time evolution.
We next consider the effect of thermal activation, which may induce tunneling processes from the external reservoirs into the SSH chain. Correspondingly, we include additional Lindblad operators of the form $\hat{L}_{i}=\hat{f}_{2i-1}^{\dagger}$ in the master equation:
[…]
where $\Gamma_{\uparrow}<\Gamma_{\downarrow}$. As shown in Fig. A7, the universal behavior remains clearly visible at a relatively large ratio $\Gamma_{\uparrow}/\Gamma_{\downarrow}=0.5$, which demonstrates the robustness of the mechanism in the presence of thermal noise. Nevertheless, since more excited states are involved in the dynamics, the entanglement values and the visibility of the revival are moderately reduced compared to the zero-temperature case, particularly at $g_{1}/g_{2}=0.1$.